# Left-Right Husimi Representation of Chaotic Resonance States: Invariance and Factorization

### v1

@@ II.1 Quantum maps
Quantum maps are well studied in quantum chaos and have a classical counterpart that is a time-discrete dynamical system [35]. For closed quantum maps the dynamics is given by a unitary time-evolution operator $U_{\text{cl}}$. We consider quantum maps with escape for which the time-evolution operator $U$ is non-unitary. If the system has a compact phase space, e.g. $\Gamma=[0,1)\times[0,1)$, $U$ can be represented by an $N\times N$ matrix, where $N$ is the dimension of the Hilbert space [36, 37]. The size $h$ of a Planck cell in phase space is then given by $h=1/N$. The time-evolution operator of the quantum map with escape is given by
@@
As the structure of resonance states depends on their decay rate $\gamma_{n}$, we also give some information about their distribution. It depends on the type of escape of the system: (i) for full escape (some columns of $U$ are zero) most decay rates are larger than the natural decay rate $\gamma_{\mathrm{nat}}$ from classical dynamics [39, 40, 41, 42, 43, 44], (ii) for partial escape most decay rates are in the interval $[\gamma_{\mathrm{nat}},\gamma_{\mathrm{inv}}]$ with $\gamma_{\mathrm{inv}}$ being the inverse natural decay rate of the time-inverted classical dynamics [38, 45, 46, 20]. In the semiclassical limit they are conjectured to accumulate around the typical decay rate $\gamma_{\mathrm{typ}}$ [38].
@@ II.2 Time-continuous scattering systems
for right resonance states $\psi^{\mathrm{R}}(\boldsymbol{r})$, where $\Delta$ is the Laplace operator in two dimensions, $n(\boldsymbol{r})$ is the local refractive index, and $k$ is the complex wavenumber. The imaginary part of $k$ is proportional to the dimensionless decay rate $\gamma$ [24, 17]. Left resonance states $\psi^{\mathrm{L}}(\boldsymbol{r})$ fulfill
$\psi^{\mathrm{L}}(\boldsymbol{r})\Delta=-n^{2}(\boldsymbol{r})k^{2}\psi^{\mathrm{L}}(\boldsymbol{r})$ (14)
with incoming boundary condition.
The LR-Husimi representation for time-continuous scattering systems is obtained by using the Husimi representation for right and left resonance states, see Eq. 12. For dielectric cavities the Husimi representation for right and left resonance states is defined in Ref. [52], specifically for right (left) resonance states we use the incoming (emanating) Husimi representation $\mathcal{H}_{n}^{\text{inc(em)}}$. For the three-disk scattering system the Husimi representation is defined in Ref. [23], where also the LR-Husimi representation is shown.
@@ II.3 Example systems
[…]
@@ III Quantum invariance
$\langle\alpha(\boldsymbol{x})|R_{n}\rangle\langle L_{n}|\alpha(\boldsymbol{x})\rangle$ =$\stackrel{{\scriptstyle(\text{i})}}{{=}}\langle\alpha(\boldsymbol{x})|U_{\text{cl}}U_{\text{cl}}^{-1}|R_{n}\rangle\langle L_{n}|U_{\text{cl}}U_{\text{cl}}^{-1}|\alpha(\boldsymbol{x})\rangle$
@@
=$\stackrel{{\scriptstyle(\text{iii})}}{{=}}\langle\alpha(T^{-1}(\boldsymbol{x}))|R_{\mathrm{refl}}|R_{n}\rangle\langle L_{n}|R_{\mathrm{refl}}^{-1}|\alpha(T^{-1}(\boldsymbol{x}))\rangle$
@@ IV.2 Factorization of LR-Husimi representation
$\mathcal{H}^{\mathrm{LR}}(\boldsymbol{x})=\frac{\sqrt{\varrho_{\gamma}^{\mathrm{L}}(\boldsymbol{x})\varrho_{\gamma}^{\mathrm{R}}(\boldsymbol{x})}\ \sqrt{\eta^{\mathrm{L}}(\boldsymbol{x})\eta^{\mathrm{R}}(\boldsymbol{x})}}{\int_{\Gamma}\text{d}\boldsymbol{x}^{\prime}\sqrt{\varrho_{\gamma}^{\mathrm{L}}(\boldsymbol{x}^{\prime})\varrho_{\gamma}^{\mathrm{R}}(\boldsymbol{x}^{\prime})}\ \sqrt{\eta^{\mathrm{L}}(\boldsymbol{x}^{\prime})\eta^{\mathrm{R}}(\boldsymbol{x}^{\prime})}}$
@@ IV.3 Distribution of LR-fluctuations
For one integration the $\delta$-function can be exploited, while the other gives an integral representation of $K_{0}$, the 00-th order modified Bessel function of the second kind [56], leading to the universal distribution of LR-fluctuations,
@@
Figure 5: Convergence of coarse-grained LR-measure $\mu_{\gamma}^{\mathrm{LR}}$ for increasingly fine partitions with $n$ cells. Shown is the Jensen-Shannon divergence $d_{\text{JS}}$ (evaluated on a $50\times 50$ grid) between measures from partitions $n$ and $n/4$ approaching zero. (a) Dielectric cavity for $\gamma\in\{0.011\,(\gamma_{\mathrm{nat}}),0.030,0.053\,(\gamma_{\mathrm{typ}}),0.090,0.122\,(\gamma_{\mathrm{inv}})\}$. (b) Three-disk scattering system for $\gamma\in\{0.436\,(\gamma_{\mathrm{nat}}),0.6,1.0,1.4,1.8\}$. (c) Standard map with partial escape for $\gamma\in\{0.22\,(\gamma_{\mathrm{nat}}),0.35,0.55,0.75,0.88\,(\gamma_{\mathrm{inv}})\}$. (d) Standard map with full escape for $\gamma\in\{0.25\,(\gamma_{\mathrm{nat}}),0.35,0.5,0.75,1.0\}$. Symbols $\circ$, $+$, $\square$, $\times$, and $\diamond$ are used for increasing $\gamma$.

### v2

@@ II.1 Quantum maps
Quantum maps are well studied in quantum chaos and have a classical counterpart that is a time-discrete dynamical system [35]. For closed quantum maps the dynamics is given by a unitary time-evolution operator $U_{\text{cl}}$. We consider quantum maps with escape for which the time-evolution operator $U$ is non-unitary. If the system has a two-dimensional compact phase space, e.g. the torus $\Gamma=[0,1)\times[0,1)$, $U$ can be represented by an $N\times N$ matrix, where $N$ is the dimension of the Hilbert space [36, 37]. The size $h$ of a Planck cell in phase space is then given by $h=1/N$. Note that this can be straightforwardly generalized to a higher-dimensional phase space. The time-evolution operator of the quantum map with escape is given by
@@
As the structure of resonance states depends on their decay rate $\gamma_{n}$, we also give some information about their distribution. It depends on the type of escape of the system: (i) for full escape (some columns of $U$ are zero) most decay rates are larger than the natural decay rate $\gamma_{\mathrm{nat}}$ from classical dynamics [39, 40, 41, 42, 43, 44], (ii) for partial escape most decay rates are in the interval $\gamma_{\mathrm{nat}},\gamma_{\mathrm{inv}}]$ with $\gamma_{\mathrm{inv}}$ being the inverse natural decay rate of the time-inverted classical dynamics [38, 45, 46, 20]. In the semiclassical limit they are conjectured to accumulate around the typical decay rate $\gamma_{\mathrm{typ}}$ [38].
@@ II.2 Time-continuous scattering systems
for right resonance states $\psi^{\mathrm{R}}(\boldsymbol{r})$, where $\Delta$ is the Laplace operator in two dimensions, $n(\boldsymbol{r})$ is the local refractive index, and $k$ is the complex wavenumber. The imaginary part of $k$ is proportional to the dimensionless decay rate $\gamma$ [24, 17]. Left resonance states $\psi^{\mathrm{L}}(\boldsymbol{r})$ fulfill an incoming boundary condition and
$\Delta\overline{\psi^{\mathrm{L}}(\boldsymbol{r})}=-n^{2}(\boldsymbol{r})\,k^{2}\,\overline{\psi^{\mathrm{L}}(\boldsymbol{r})}\,.$ (14)
The LR-Husimi representation for time-continuous scattering systems is obtained by using the Husimi representation for right and left resonance states, see Eq. 12. For dielectric cavities the Husimi representation for right and left resonance states is defined in Ref. [52]. Specifically, for right and left resonance states we use the incoming Husimi representation $\mathcal{H}_{n}^{\text{inc}}$ inside the cavity. Alternatively, one could use for both states the emanating Husimi representation, giving the same LR-Husimi representation in the semiclassical limit. In this context the terms incoming and emanating refer to the definition of the boundary Husimi representation inside the cavity in Ref. [52] and are not related to the incoming and outgoing boundary conditions of left and right resonance states. For the three-disk scattering system the Husimi representation is defined in Ref. [23], where also the LR-Husimi representation is shown.
@@ II.3 Example systems
In all figures, for the most semiclassical resonance states we use the following ranges of wavenumbers: (a) for the dielectric cavity $\mathrm{Re}\,kR_{\mathrm{cav}}\in[5\,000,5\,025]$, (b) for the three-disk scattering system $\mathrm{Re}\,ka\in[100\,000,100\,200]$ (except for $\gamma=1.8$ in Figs. 6, 8, and 10, where $\mathrm{Re}\,ka\in[50\,000,50\,500]$ is used), and (c, d) for the standard map with partial and full escape all even $N\in[19\,996,20\,004]$.
We expect that our results hold for higher-dimensional billiards and maps with fully chaotic phase spaces just as well, since all our arguments are independent of the number of dimensions of phase space. However, those systems are harder to visualize and numerically more challenging, in particular they would not allow us to go as far into the semiclassical limit. Therefore, we focus on the lowest dimensional billiards and maps with chaotic dynamics.
@@ III Quantum invariance
$\langle\alpha(\boldsymbol{x})|R_{n}\rangle\langle L_{n}|\alpha(\boldsymbol{x})\rangle$ $\stackrel{{\scriptstyle(\text{i})}}{{=}}\langle\alpha(\boldsymbol{x})|U_{\text{cl}}U_{\text{cl}}^{-1}|R_{n}\rangle\langle L_{n}|U_{\text{cl}}U_{\text{cl}}^{-1}|\alpha(\boldsymbol{x})\rangle$
@@
$\stackrel{{\scriptstyle(\text{iii})}}{{=}}\langle\alpha(T^{-1}(\boldsymbol{x}))|R_{\mathrm{refl}}|R_{n}\rangle\langle L_{n}|R_{\mathrm{refl}}^{-1}|\alpha(T^{-1}(\boldsymbol{x}))\rangle$
@@ IV.2 Factorization of LR-Husimi representation
$\mathcal{H}^{\mathrm{LR}}(\boldsymbol{x})=\frac{\sqrt{\varrho_{\gamma}^{\mathrm{L}}(\boldsymbol{x})\varrho_{\gamma}^{\mathrm{R}}(\boldsymbol{x})}\ \sqrt{\eta^{\mathrm{L}}(\boldsymbol{x})\eta^{\mathrm{R}}(\boldsymbol{x})}}{\int_{\Gamma}\text{d}\boldsymbol{x}^{\prime}\sqrt{\varrho_{\gamma}^{\mathrm{L}}(\boldsymbol{x}^{\prime})\varrho_{\gamma}^{\mathrm{R}}(\boldsymbol{x}^{\prime})}\ \sqrt{\eta^{\mathrm{L}}(\boldsymbol{x}^{\prime})\eta^{\mathrm{R}}(\boldsymbol{x}^{\prime})}}$ (23)
@@ IV.3 Distribution of LR-fluctuations
For one integration the $\delta$-function can be exploited, while the other gives an integral representation of $K_{0}$, the 0-th order modified Bessel function of the second kind [56], leading to the universal distribution of LR-fluctuations,
@@
Figure 5: Convergence of coarse-grained LR-measure $\mu_{\gamma}^{\mathrm{LR}}$ for increasingly fine partitions with $n$ cells. Shown is the Jensen-Shannon divergence $d_{\text{JS}}$ (evaluated on a $50\times 50$ grid) between measures from partitions $n$ and $n/4$ approaching zero. (a) Dielectric cavity for $\gamma\in\{0.011\,(\gamma_{\mathrm{nat}}),0.030,0.053\,(\gamma_{\mathrm{typ}}),0.090,0.122\,(\gamma_{\mathrm{inv}})\}$. (b) Three-disk scattering system for $\gamma\in\{0.436\,(\gamma_{\mathrm{nat}}),0.6,1.0,1.4,1.8\}$. (c) Standard map with partial escape for $\gamma\in\{0.22\,(\gamma_{\mathrm{nat}}),0.35,0.55,0.75,0.88\,(\gamma_{\mathrm{inv}})\}$. (d) Standard map with full escape for $\gamma\in\{0.25\,(\gamma_{\mathrm{nat}}),0.35,0.5,0.75,1.0\}$. Symbols $\circ$, ++, $\square$, $\times$, and $\diamond$ are used for increasing $\gamma$.
@@ VI Summary and Outlook
An important open question is how the presented results can be observed in experimental scattering systems. As we are in a regime of strongly overlapping resonances, the properties of an individual resonance are experimentally not accessible. However, the average over many resonances with similar decay rates also depends on the decay rate and might be observable by the time evolution of wave packets. This is experimentally related to pulse propagation deduced from transmission spectra [62].
Data availability statement
The data that support the findings of this study are openly available at the following URL/DOI: https://doi.org/10.25532/OPARA-989.